Refer to the data set in the accompanying table. Assume that the paired sample data is a simple random sample and the differences have a distribution that is approximately normal. Use a significance level of 0.01 to test for a difference between the weights of discarded paper? (in pounds) and weights of discarded plastic? (in pounds).

Household Paper Plastic
1 9.41 3.36
2 15.09 9.11
3 6.83 3.57
4 7.98 6.09
5 17.65 11.26
6 7.57 5.92
7 13.31 19.7
8 9.19 3.74
9 12.32 11.17
10 13.61 8.95
11 16.08 14.36
12 6.44 8.4
13 11.08 12.47
14 9.55 9.2
15 2.8 5.92
16 11.42 12.81
17 20.12 18.35
18 9.45 3.02
19 6.38 8.82
20 6.98 2.65
21 13.05 12.31
22 16.39 9.7
23 6.96 7.6
24 6.67 6.09
25 5.86 3.91
26 7.72 3.86
27 8.72 9.2
28 11.36 10.25
29 6.16 5.88
30 6.33 3.86

t=?

p=?
Answer #1

x1=paper , x2 = plastic

First enter Data into EXCEL

We have to find the sample mean.

Excel command is =AVERAGE(Select data)

sample mean =\bar{x1} = 10.08

sample mean =\bar{x2} = 8.38

Now we have to find sample standard deviation.

Excel command is =STDEV(Select data)

standard deviation = s1 = 4.11

standard deviation = s2 = 4.40

n1 = 30

n2 = 30

\bar{x1}= 10.08

\bar{x2} = 8.38

s1 = 4.11

s2 = 4.40

Null and alternative hypothesis is

H0 :u1 = u2

Vs

H1 :u1? u2

Level of significance = 0.01

Before doing this test we have to check population variances are equal or not.

Null and alternative hypothesis is

H0:\sigma _{1}^{2}=\sigma _{2}^{2}

Vs

H1 : :\sigma _{1}^{2 } ? \sigma _{2}^{2} ? \sigma _{2}^{2}

Test statistic is

F = Larger variance / Smaller variance = 19.36 / 16.8921 = 1.146

Degrees of freedoms

Degrees of freedom for numerator = n1 - 1 = 30 - 1 = 29

Degrees of freedom for denominator = n2 - 1 = 30 - 1 = 29

Critical value = 2.42            ( using f-table )

F test statistic < critical value    we fail to reject null hypothesis.

Conclusion: Population variances are equal.

So we have to use pooled variance.

Formula

t=\frac{\bar{x1}-\bar{x2}}{Sp*\sqrt{\frac{1}{n1}+\frac{1}{n2}}}

Sp = \sqrt{\frac{(n1-1)s1^{2}+(n2-1)s2^{2}}{n1+n2-2}}

30-1)16.8921+(30-1)19.36 30+30-2

Sp = 4.257

t=\frac{10.08 - 8.38 }{ 4.257 *\sqrt{\frac{1}{30}+\frac{1}{30}}}

t=1.546

? = 0.01 , d.f = n1 + n2 – 2 = 3 0+ 30 - 2 = 58

p-value = 0.127 ( using t table )

p-value > ? ,Failed to Reject H0